Solve the equation 𝑆 = ∫10𝑒2t 𝑑𝑑 for T range 0 to T.




1
Expert's answer
2021-12-02T15:17:49-0500

Solution;

"S=\displaystyle \int_0^T10e^{2t}dt"


"S=10\displaystyle \int_0^Te^{2t}dt"


"S=10[\frac{e^{2t}}{2}]_0^T"

"S=5(e^{2T}-e^{0})"

"S=5(e^{2T}-1)"